8.1. To show xn 2 Xn+1 for n 2 N and to find the value of N EN, consider the function. f: (0, +00) → R and so f(t) = n) ObserIf you could please answer 8.1 & 8.2 that would be incredible. Thank you so much.
Answer #1

Senti = 3 ton) Then S2 = 3 Ctsi) = 7CH) z = f (H2) = 3 (1H2) = 1/2 Sy = $(133) = $Cho 55 = $(457) = $(14) ) = ③ we pore sn= 2© Brom @ clesly six Sz 6;1>$) The result is tree fal nal Assume that result is there for n=k Them Sk> Sato Phat is ShHTSA Sitfrom Snte = 11t5n). hormon that a home (145) d = 3 Citd) 3d=HU zdal d=/ Honce imat